Estimate the firm's stock price using the information below
Estimated FCFF 14 million every year, forever
WACC 12.1%
Short-term investments 2 million
Debt 32 million
Preferred stock 8 million
Shares of common stock 5.1 million
SHOW WORK HERE, HIGHLIGHT FINAL ANSWER IN YELLOW
Estimate the firm's stock price using the information below
Current period dividend 8.90 dollars per share
Required return on equity 20%
Constant growth rate (g) 2%
Shares of common stock 122 million
SHOW WORK HERE, HIGHLIGHT FINAL ANSWER IN YELLOW
Answer #1

1). Firm Value = Expected FCF / WACC = 14 / 0.121 = $115.70 million

Equity Value = Firm Value - Debt - Preferred Stock + Short-term investments

= 115.70 - 32 - 8 + 2 = $77.70 million

Share Price = Equity Value / No. of outstanding value = 77.70 / 5.1 = $15.24

2). Share Price = [Current Dividend * (1 + g)] / [r - g]

= [$8.90 * (1 + 0.02)] / [0.20 - 0.02] = $9.078 / 0.18 = $50.43